OpenStudy (anonymous):

find 2 values of x for where f(x)=0. f(x) = −2 cos(2x + π) + 1

myininaya (myininaya):

You are being asked to solve -2cos(2x+pi)+1=0

myininaya (myininaya):

isolate the trig function part.

myininaya (myininaya):

Then use the handy dandy unit circle.

OpenStudy (dan815):

what makes it so handy dandy

OpenStudy (anonymous):

I don't understand... so like cos(2x+pi) = 0?

OpenStudy (ranga):

-2cos(2x+pi)+1=0 add +2cos(2x+pi) to both sides 0 + 1 = 2cos(2x+pi) 2cos(2x+pi) = 1 divide both sides by 2 cos(2x+pi) = 1/2 Find two values when cosine is 1/2. Substitute and solve for x.

OpenStudy (anonymous):

2 values when cosine is 1/2 where? If I graph it would it help?

OpenStudy (anonymous):

do you mean on the unit circle? because then I believe that is pi/3 and 5pi/3

OpenStudy (ranga):

calculator will tell you. It is called arc cos or inverse cosine marked with cos^-1. Or you can look up trig tables.

OpenStudy (ranga):

yes pi/3 and 5pi/3 are good. 2x + pi = pi/3. solve for x. 2x + pi = 5pi/3. solve for x.

OpenStudy (anonymous):

what about a value for x when f(x) is minimized? any clue what that means?

OpenStudy (ranga):

cosine function goes from +1 to -1. f(x) = −2 cos(2x + π) + 1 Put the max and min values for cosine and find f(x) f(x) = -2(1) + 1 = -1 f(x) = -2(-1) + 1 = 3 So f(x) has a minimum when the cosine function is maximum at 1. Solve for x when cos(2x + pi) = 1

OpenStudy (anonymous):

any suggestions about finding the range of f(x)?

OpenStudy (anonymous):

and is the answer to the minimizing just 2x+pi=2pi and solve? because thats the only time cosine = 1, right?